OpenStudy (anonymous):

the sum of 4 times a number is 14 is 16. What is the number?

OpenStudy (accessdenied):

I think you mean "the sum of ... AND 14 is 16", in which case... We'll call the number 'x' 4*x + 14 = 16 4x = 2 x = 2/4 = 1/2